# Conservation laws in non-inertial frames and non-conservation of energy of relative motion in two-body problem

### v2

@@ Abstract
The dynamics of systems of multiple gravitationally interacting bodies is often studied in a frame attached to one of the objects (e.g. a central star in a planetary system). As this frame is generally non-inertial, indirect forces appear in the equations describing the motion of bodies relative to the reference object. According to the convention adopted in celestial mechanics, the associated indirect acceleration is defined to be different for every object under consideration, whereas the gravitational coupling between each body and the reference object is described via the effective two-body potential, which does not obey the equivalence principle. Here we point out that a slightly different and more physically motivated definition of the indirect acceleration provides significant benefits when interpreting relative motion in a non-inertial frame. First, the indirect acceleration ends up being the same for all objects in the system. Second, the non-conservation of momentum, angular momentum, and energy of the whole system in a non-inertial frame naturally follow from the action of the indirect acceleration on the system as an external force. We also argue that the vis viva integral of the classical two-body problem should not be interpreted as a statement of energy conservation. In fact, the energy of relative motion is not conserved due to the work done on the two-body system by the indirect force. These results can be useful for interpreting dynamics in various astrophysical contexts, in particular the physics of disc-planet coupling.
@@ 1 Introduction
While mathematical implications of the presence of indirect forces are well understood in celestial mechanics, physical interpretation of these forces in their conventional form can be tricky. In this pedagogical note, we propose a reformulation of the indirect forces (Section 3), which is slightly different from the classical approach (recited in Section 2) used to describe dynamics in a non-inertial frame centered on one of the bodies. We show that this reformulation leads, in particular, to a very straightforward interpretation of conservation laws in a non-inertial system (Section 4). We also re-examine the classical two-body problem (Section 5) and demonstrate that some of its well-established aspects, e.g. the vis viva integral and its relation to energy conservation, take on a new meaning in light of the proposed reformulation. Mathematical details of our derivations can be found in the Appendix.
@@ 2 Dynamics in a non-inertial frame and indirect forces: conventional view
Then, using equation (1), we can write down the equation of motion for $m_{n}$ in the frame co-moving with $m_{1}$ as (e.g. Tremaine2023)
@@
The decomposition given by equations (4)-(10) of $\ddot{\bf R}_{n}$ into different contributions follows a standard convention adopted in celestial mechanics (e.g. Smart1953; Brou1961; Murray1999; Tremaine2023). However, it is not ideal for a number of reasons. First, the indirect acceleration represented by the last sum in equation (3) excludes the gravitational effect of mass $m_{n}$ on $m_{1}$, and there is no a priori physical reason why this should be the case. Second, let us for a moment consider the two-body problem, when $N=2$. Then one finds $\Phi_{\mathrm{d}}=\tilde{\Phi}_{\mathrm{id}}=0$ so that the equation (3) for the motion of mass $m_{2}$ reduces to a classical result
@@
where ${\bf R}_{2}={\bf r}_{2}-{\bf r}_{1}$. The gravitational coupling responsible for the evolution of ${\bf R}_{2}$ is thus represented entirely by the effective two-body potential $\Phi_{\mathrm{2b}}=-G(m_{1}+m_{2})/R_{2}$. However, $\Phi_{\mathrm{2b}}$ is not a true gravitational potential as it does not satisfy the equivalence principle: according to equation (11), the acceleration of a mass $m_{2}$ due to $m_{1}$ depends on the mass $m_{2}$ itself, which violates the equivalence principle. Same issue arises in the general $N$-body case, see equation (8). We will discuss two-body problem in greater depth in Section 5.
@@ 3 An alternative interpretation of the dynamics in a non-inertial frame
This equation enables a very straightforward interpretation of the force balance in a non-inertial frame of $m_{1}$: accelerated motion of $m_{1}$ causes the emergence of an acceleration $\textbf{{a}}_{\mathrm{id}}$ in the frame co-moving with $m_{1}$, and this (reflex) acceleration acts uniformly on all other $n=2,\dots,N$ objects in the system (see also Crida2025). As a result, the acceleration $\ddot{\bf R}_{n}$ experienced by $m_{n}$ is contributed to by the classical gravitational attraction of $m_{n}$ to all other objects in the system (second term $\textbf{{a}}_{\mathrm{d}}$), including $m_{1}$ (first term $\textbf{{a}}_{\mathrm{N}}$), as well the reflex acceleration $\textbf{{a}}_{\mathrm{id}}$ (the same for all $n=2,\dots,N$ objects) due to the non-inertial motion of $m_{1}$. This simple physical interpretation is lacking in the conventional approach described in Section 2.
@@
In such cases, when studying the dynamics of mass $m_{n}$, in a zeroth-order approximation one would set all $m_{k}=0$, $k=2,\dots,N$, in the equation (12). In this limit $\textbf{{a}}_{\mathrm{d}}=\textbf{{a}}_{\mathrm{id}}={\bf 0}$, so that equation (15) reduces to equation $\ddot{\bf R}_{n}=-(Gm_{1}/R_{n}^{3}){\bf R}_{n}$ describing relative motion of a test mass in the classical two-body problem. The neglected terms proportional to $m_{k}$, $k=2,\dots,N$, get added at the next level of approximation as a perturbation, as is normally done in celestial mechanics.
@@ 5 Two-body problem
The vis viva integral is commonly interpreted as the ‘energy integral’ of motion in the potential $\Phi_{\mathrm{2b}}$, implying that the orbital energy per unit mass is conserved (Murray1999; Kop2011). However, we find this interpretation rather misleading.
First, as we argued in Section 2, $\Phi_{\mathrm{2b}}$ does not represent a true gravitational potential of the system and raises issues with the equivalence principle. Second, the extension of the definition of energy $E$ to the case of $N>2$ based on the $N=2$ expression for $I$ (i.e. interpreting $m_{n}\Phi_{\mathrm{2b}}$ as a potential energy of gravitational coupling with $m_{1}$ instead of $m_{n}\Phi_{\mathrm{N}}$) is not obvious, cf. equation (22). Third, equation (29) simply does not represent the total energy of the relative motion of the two-body system, i.e. the motion of $m_{2}$ as viewed from a frame co-moving with $m_{1}$. The true energy $E$ — the sum of kinetic and potential energies in the frame of $m_{1}$— is given by the equation (24) for $N=2$, namely
$E=\frac{1}{2}m_{2}\dot{\bf R}_{2}^{2}-\frac{Gm_{1}m_{2}}{R_{2}}\,.$ (30)
It is clear that $E$ is different from $I$, even when the latter is multiplied by $m_{2}$. But this is hardly surprising if one resorts to our interpretation of the relative motion advanced in Section 3.
@@
so that the energy of the relative motion $E$ is not an integral of motion. This equation reflects the non-conservation of $E$ due to the work done on the two-body system by the indirect force.
@@
where we used equations (29) and (30). In other words, the IoM $m_{2}I$ is only the ‘energy-like’ IoM that represents the conservation of the sum of the total energy of the system $E$ and the work done upon it by the indirect force $M\textbf{{a}}_{\mathrm{id}}$. This analysis illustrates why calling $I$ (or $m_{2}I$) the ‘energy integral’ can be misleading. Using this specific name in relation to $I$ should be avoided.
@@ 6 Discussion
The main goal of this work was to highlight the benefits of using the modified form (14) of the indirect acceleration $\textbf{{a}}_{\mathrm{id}}$ in comparison with the standard form (10) conventionally used in celestial mechanics. It should be emphasized that using this form of $\textbf{{a}}_{\mathrm{id}}$ does not in any way change the actual calculations of motion of objects in $N$-body systems. The laws of motion given by equations (3) and (15) are identical to each other, regardless of how we partition their different terms.
@@
This interpretation is to be contrasted with e.g. the classical ‘conservation law’ (29) in a two-body system, that requires (i) the introduction of a special 2-body potential $\Phi_{\mathrm{2b}}$ that does not obey the equivalence principle and (ii) the definition of ‘energy’ $I$ that cannot be easily extended to systems of $N>2$ objects.
Equations (26)-(28) illustrate the instantaneous non-conservation of $\bf P$}, $\bf L$}, and $E$. However, over timescales long compared to the dynamical timescale of the system, their right-hand sides should average out to zero, leading to the usual conservation of the total energy and angular momentum of systems of planets in a secular approximation (at least as long as these systems remain finite and bound). In fact, indirect forces are known to have no effect in secular perturbation theory in celestial mechanics (Murray1999).
Interestingly, standard tests of the performance of various integration algorithms in celestial mechanics are almost exclusively focused on integrating the two-body problem and observing the conservation of the vis-viva integral $I$, instead of the energy $E$. As a result, they would miss the short-term energy non-conservation described by equation (28) in more complicated setups.
@@ 7 Summary
In this approach the indirect acceleration ends up being the same for all objects of the system (which is different from the standard convention adopted in celestial mechanics) and affects the system’s dynamics as an external force, leading to instantaneous non-conservation of $\bf P$}, $\bf L$}, and $E$. This non-conservation has important implications not only for $N$-body systems but also for disc-planet interaction — gravitational coupling between a massive planet and a fluid disc, as we show in Rafikov2025.
We also demonstrate that the vis-viva integral of the two-body problem, which is classically considered to reflect energy conservation for the relative motion of two gravitating masses, describes energy non-conservation in our interpretation: the total energy associated with the relative motion of two gravitating objects changes due to the work done by the indirect force. Thus, the vis-viva integral is only an energy-like integral of motion, and not the true energy (sum of kinetic and potential energies) of relative motion in a two-body system.
@@ Acknowledgements
The author is grateful to Callum Fairbairn and Alexander Dittmann for their comments on the manuscript, and to Scott Tremaine and Nicolas Cimerman for useful discussions. R.R.R. acknowledges financial support from the STFC grant ST/T00049X/1 and the IBM Einstein Fellowship at the IAS.

### v3

@@ Abstract
The dynamics of systems of multiple gravitationally interacting bodies is often studied in a frame attached to one of the objects (e.g. a central star in a planetary system). As this frame is generally non-inertial, indirect forces appear in the equations describing the motion of bodies relative to the reference object. According to the convention adopted in celestial mechanics, the associated indirect acceleration is defined to be different for every object under consideration, whereas the gravitational coupling between each body and the reference object is described via the effective two-body potential, which does not obey the equivalence principle. Here we point out that a slightly different and more physically motivated definition of the indirect acceleration provides significant benefits when interpreting relative motion in a non-inertial frame. First, the indirect acceleration ends up being the same for all objects in the system. Second, the non-conservation of momentum, angular momentum, and energy of the whole system in a non-inertial frame naturally follow from the action of the indirect acceleration on the system as an external force. We also argue that the vis viva integral of the classical two-body problem should not be interpreted as a statement of energy conservation in a non-inertial frame attached to one of the bodies. The energy of relative motion is not conserved in this frame due to the work done on the two-body system by the indirect force. These results can be useful for interpreting dynamics in various astrophysical contexts, in particular the physics of disc-planet coupling.
@@ 1 Introduction
While mathematical implications of the presence of indirect forces are well understood in celestial mechanics, physical interpretation of these forces in their conventional form can be tricky. In this pedagogical note, we propose a reformulation of the indirect forces (Section 3), which is slightly different from the classical approach (recited in Section 2) used to describe dynamics in a non-inertial frame centered on one of the bodies. We show that this reformulation leads, in particular, to a very straightforward interpretation of conservation laws in a non-inertial system (Section 4). This interpretation is one of the key motivations for the present work and will be used in a companion study (Rafikov et al., 2025) of disc-planet interaction in a non-inertial frame. We also re-examine the classical two-body problem (Section 5) and demonstrate that some of its well-established aspects, e.g. the vis viva integral and its relation to energy conservation, take on a new meaning in light of the proposed reformulation. Mathematical details of our derivations can be found in the Appendix.
@@ 2 Dynamics in a non-inertial frame and indirect forces: conventional view
Then, using equation (1), we can write down the equation of motion for $m_{n}$ in the frame co-moving with $m_{1}$ as (e.g. Tremaine, 2023)
@@
The decomposition given by equations (4)-(10) of $\ddot{\bf R}_{n}$ into different contributions follows a standard convention adopted in celestial mechanics (e.g. Smart, 1953; Brouwer and Clemence, 1961; Murray and Dermott, 1999; Tremaine, 2023). However, it is not ideal for a number of reasons. First, the indirect acceleration represented by the last sum in equation (3) excludes the gravitational effect of mass $m_{n}$ on $m_{1}$, and there is no a priori physical reason why this should be the case. Second, let us for a moment consider the two-body problem, when $N=2$. Then one finds $\Phi_{\mathrm{d}}=\tilde{\Phi}_{\mathrm{id}}=0$ so that the equation (3) for the motion of mass $m_{2}$ reduces to a classical result
@@
where ${\bf R}_{2}={\bf r}_{2}-{\bf r}_{1}$. In the classical interpretation, the effective two-body potential $\Phi_{\mathrm{2b}}=-G(m_{1}+m_{2})/R_{2}$ fully represents the gravitational coupling responsible for the evolution of ${\bf R}_{2}$, with no other forces (except gravity) present in the two-body system. However, $\Phi_{\mathrm{2b}}$ is not a true gravitational potential as it does not satisfy the equivalence principle: according to equation (11), the acceleration of a mass $m_{2}$ due to $m_{1}$ depends on the mass $m_{2}$ itself, which violates the equivalence principle. Same issue arises in the general $N$-body case, see equation (8). We will discuss two-body problem in greater depth in Section 5.
@@ 3 An alternative interpretation of the dynamics in a non-inertial frame
This equation enables a very straightforward interpretation of the force balance in a non-inertial frame of $m_{1}$: accelerated motion of $m_{1}$ causes the emergence of an acceleration $\textbf{{a}}_{\mathrm{id}}$ in the frame co-moving with $m_{1}$, and this (reflex) acceleration acts uniformly on all other $n=2,\dots,N$ objects in the system (see also Crida et al., 2025). As a result, the acceleration $\ddot{\bf R}_{n}$ experienced by $m_{n}$ is contributed to by the classical gravitational attraction of $m_{n}$ to all other objects in the system (second term $\textbf{{a}}_{\mathrm{d}}$), including $m_{1}$ (first term $\textbf{{a}}_{\mathrm{N}}$), as well the reflex acceleration $\textbf{{a}}_{\mathrm{id}}$ (the same for all $n=2,\dots,N$ objects) due to the non-inertial motion of $m_{1}$. This simple physical interpretation is lacking in the conventional approach described in Section 2.
@@
In such cases, when studying the dynamics of mass $m_{n}$, in a zeroth-order approximation one would set all $m_{k}=0$, $k=2,\dots,N$, in the equation (12). In this limit $\textbf{{a}}_{\mathrm{d}}=\textbf{{a}}_{\mathrm{id}}={\bf 0}$, so that equation (15) reduces to equation $\ddot{\bf R}_{k}=-(Gm_{1}/R_{k}^{3}){\bf R}_{k}$ describing relative motion of a test mass in the classical two-body problem. The neglected terms proportional to $m_{k}$, $k=2,\dots,N$, would get added at the next level of approximation as a perturbation, as is normally done in celestial mechanics.
Note that this procedure does not provide an exact solution of the two-body problem for each individual object $m_{k}$ ($k>1$) at the zeroth-order approximation. This is because the angular frequency of the test mass problem $n^{2}=Gm_{1}/a_{k}^{3}$ ($a_{k}$ is the semi-major axis of $m_{k}$) is slightly different from the two-body problem with a finite $m_{k}$, $n^{2}=G(m_{1}+m_{k})/a_{k}^{3}$. To remedy this problem, one can employ a staged approach to adding masses $m_{i}$ into the consideration: for every mass $m_{k}$ one would first add only the terms involving $m_{k}$, keeping all other $m_{i}=0$ ($i>1,i\neq k$); at this level of approximation $\Phi_{\mathrm{d}}=0$ while $\Phi_{\mathrm{id}}=-Gm_{k}/R_{k}$, and the classical expression for the angular frequency of the two-body problem is easily recovered. Once this is done, at the next level of approximation, one would include all other terms proportional to $m_{i}=0$ ($i>1,i\neq k$).
@@ 5 Two-body problem
The vis viva integral is commonly interpreted as the ‘energy integral’ of motion in the potential $\Phi_{\mathrm{2b}}$, implying that the orbital energy per unit mass is conserved (Murray and Dermott, 1999; Kopeikin et al., 2011). Indeed, one can show that $\mu I$, where $\mu=m_{1}m_{2}/(m_{1}+m_{2})$ is the reduced mass, represents the energy of a two-body system computed in the inertial, barycentric frame (Murray and Dermott, 1999; Tremaine, 2023).
However, the vis viva integral is often derived by considering relative motion of two objects in a non-inertial frame attached to one of them (Murray and Dermott, 1999). In that frame the interpretation of $I$ as the ’energy integral’ is rather misleading. First, as we argued in Section 2, $\Phi_{\mathrm{2b}}$ does not represent a true gravitational potential of the system and raises issues with the equivalence principle. It should not be considered as the gravitational potential energy of the system.
Second, the extension of the definition of energy $E$ to the case of $N>2$ based on the $N=2$ expression for $I$ (i.e. interpreting $m_{n}\Phi_{\mathrm{2b}}$ as a potential energy of gravitational coupling with the mass $m_{1}$ instead of the true potential energy $m_{n}\Phi_{\mathrm{N}}$) is not obvious. Indeed, it does not seem possible to formulate a physically-motivated analog of equation (22) that would involve $\Phi_{\mathrm{2b}}$.
Third, equation (29) simply does not represent the total energy of the relative motion of the two-body system, i.e. the motion of $m_{2}$ as viewed from a frame co-moving with $m_{1}$. The true energy $E$ — the sum of kinetic and gravitational potential energies in the frame of $m_{1}$— is given by the equation (24) for $N=2$, namely
$E=\frac{1}{2}m_{2}\dot{\bf R}_{2}^{2}+m_{2}\Phi_{\mathrm{N}}=\frac{1}{2}m_{2}\dot{\bf R}_{2}^{2}-\frac{Gm_{1}m_{2}}{R_{2}}\,.$ (30)
It is clear that $E$ is different from $I$, even when the latter is multiplied by $m_{2}$. But this is hardly surprising if one resorts to our interpretation of the relative motion in a non-inertial frame advanced in Section 3.
@@
so that the energy of the relative motion $E$ is not an integral of motion. This equation reflects the non-conservation of $E$ due to the work done on the two-body system by the indirect force. Note that in our interpretation $E$ is defined with the true gravitational potential $\Phi_{\mathrm{N}}$ satisfying the equivalence principle, but unlike the classical two-body interpretation (see Section 2), there is an additional force in the system — the indirect force.
@@
where we used equations (29) and (30). In other words, in a non-inertial frame the IoM $m_{2}I$ is only the ‘energy-like’ IoM that represents the conservation of the sum of the total energy of the system $E$ and the work done upon it by the indirect force $M\textbf{{a}}_{\mathrm{id}}$. This analysis illustrates why calling $I$ (or $m_{2}I$) the ‘energy integral’ can be misleading when considering motion in the frame of one of the bodies. Using this specific name in relation to $I$ in that frame should be avoided.
@@ 6 Discussion
One of the main goals of this work was to highlight the benefits of using the modified form (14) of the indirect acceleration $\textbf{{a}}_{\mathrm{id}}$ in comparison with the standard form (10) conventionally used in celestial mechanics. It should be emphasized that using this form of $\textbf{{a}}_{\mathrm{id}}$ does not in any way change the actual calculations of motion of objects in $N$-body systems. The laws of motion given by equations (3) and (15) are identical to each other, regardless of how we partition their different terms.
@@
This interpretation is to be contrasted with e.g. the classical ‘conservation law’ (29) for a two-body system in a non-inertial frame, that requires (i) the introduction of a special 2-body potential $\Phi_{\mathrm{2b}}$ that does not obey the equivalence principle and (ii) the definition of ‘energy’ $I$ that cannot be easily extended to systems of $N>2$ objects.
Equations (26)-(28) illustrate the instantaneous non-conservation of $\bf P$}, $\bf L$}, and $E$ in a non-inertial frame. However, over timescales long compared to the dynamical timescale of the system, their right-hand sides should average out to zero, leading to the usual conservation of the total energy and angular momentum of systems of planets in a secular approximation (at least as long as these systems remain finite and bound). In fact, indirect forces are known to have no effect in secular perturbation theory in celestial mechanics (Murray and Dermott, 1999).
Interestingly, standard tests of the performance of various integration algorithms in celestial mechanics are often focused on integrating the two-body problem and observing the conservation of energy in an inertial, barycentric frame (which is the same as the vis-viva integral $I$, up to a constant factor). Since this frame is different from a non-inertial frame attached to one of the objects, the short-term energy non-conservation described by equation (28) is not observed in such tests.
Understanding indirect forces is important not only in celestial mechanics. For example, they also arise in studies of disc-planet interaction Goldreich and Tremaine (1980), for which a natural coordinate system is the one attached to the central star. Crida et al. (2025) and Rafikov et al. (2025) explored the role played by the indirect force in disc-planet studies, finding it to be important in situations where the planet is massive enough to significantly perturb the disc and open a gap around its orbit. Rafikov et al. (2025) also examined global angular momentum balance of the disc-planet system in the non-inertial stellar frame, showing that the full angular momentum of the system is not conserved. This finding is fully in line with our equation (27) demonstrating non-conservation of $\bf L$} in non-inertial frames.
@@ 7 Summary
In this approach the indirect acceleration ends up being the same for all objects of the system (which is different from the standard convention adopted in celestial mechanics) and affects the system’s dynamics as an external force, leading to instantaneous non-conservation of $\bf P$}, $\bf L$}, and $E$. This non-conservation has important implications not only for $N$-body systems but also for disc-planet interaction — gravitational coupling between a massive planet and a fluid disc, as we show in Rafikov et al. (2025).
We also demonstrate that the vis-viva integral of the two-body problem, which is classically considered to reflect energy conservation for the relative motion of two gravitating masses, in our interpretation describes energy non-conservation in a non-inertial frame: the total energy associated with the relative motion of two gravitating objects in that frame changes due to the work done by the indirect force. Thus, the vis-viva integral is only an energy-like integral of motion in a non-inertial frame attached to one of the objects, and not the true energy (sum of kinetic and potential energies) of relative motion in a two-body system.
@@ Acknowledgements
The author is grateful to Callum Fairbairn and Alexander Dittmann for their comments on the manuscript, to Scott Tremaine and Nicolas Cimerman for useful discussions, and to Daniel Tamayo and an anonymous referee for their suggestions that helped sharpen the logic of the manuscript. R.R.R. acknowledges financial support from the STFC grant ST/T00049X/1 and the IBM Einstein Fellowship at the IAS.
Data availability
No new data were generated or analysed in support of this research.